OpenStudy (anonymous):

(square root 2 sin (x))+1+=0 Solve the equation for x. Use n as an integer constant.

OpenStudy (whpalmer4):

What do you want to do with it?

OpenStudy (anonymous):

Solve the equation for x. Use n as an integer constant.

OpenStudy (anonymous):

can you show me step by step how to do the problem

OpenStudy (anonymous):

\[\sqrt{2}\sin x+1=0?\\ \sqrt{2}\sin x=-1\\ \sin x=-\frac{1}{\sqrt2}\] What values of x make this true?

OpenStudy (anonymous):

I don't know...:(

OpenStudy (anonymous):



OpenStudy (anonymous):

I am appreciating you trying to help but I have no idea what that picture is...I thought I was suppose to looking on the unit circle at some point..?

OpenStudy (anonymous):

 Pretend that's a circle. The triangles are reference triangles containing the angle x in the equation. The outermost vertex of each triangle lies on the unit circle.